OpenStudy (anonymous):

What is the general form of the equation of a circle with center at (a, b) and radius of length m?

OpenStudy (anonymous):

x2 + y2 − 2ax − 2by + (a2 + b2 − m2) = 0 x2 + y2 + 2ax + 2by + (a2 + b2 − m2) = 0 x2 + y2 − 2ax − 2by + (a + b − m2) = 0 x2 + y2 + 2ax + 2by + a2 + b2 = -m2 plz explain

OpenStudy (mathstudent55):

The equation of a circle with center (a, b) and radius m is: \((x - a)^2 + (y - b)^2 = m^2\) Now take that equation, square both binomials and simplify. Then match what you get to one of the choices.

OpenStudy (mathstudent55):

\((x - a)^2 + (y - b)^2 = m^2\) \(x^2 - 2ax + a^2 + y^2 - 2by + b^2 = m^2\) \(x^2 + y^2 - 2ax - 2by + a^2 + b^2 - m^2 = 0\) \(x^2 + y^2 - 2ax - 2by + (a^2 + b^2 - m^2) = 0\)